1. If A= [ 1 = [] 4]. find the general solution to dx dt Ax
Answer #1

Given, A= [ ] To find the General Solution of ht Eigenvalues : =Ax A-XI= 2-1 - = O - 1 4-X = (2-1) (4-1) + 1 = 0 or 8 - 27-44Second e linearly independent solution : - Now, we need to solve A (second vector) -1 (second veeter) = (flest vector) or [ ]